OpenStudy (anonymous):

Solve for x x^2-10x+106=0

OpenStudy (anonymous):

will x be imaginary???

OpenStudy (anonymous):

yes.. but i dont know how to figure it out.. can you help please

jhonyy9 (jhonyy9):

D=100-424=-324 x_1,2=(10 +/- isqrt324)/2 = (10 +/- isqrt(4*81))/2 = (10+/- 18i)/2 = = 5+/- 9i

OpenStudy (anonymous):

thanks